UARS Weekly Status Report

04 May 2001




********************************************************************

                UARS WEEKLY STATUS REPORT - 4 May 2001

********************************************************************

***********
SOLSTICE
***********

One operational anomaly occured this week that required SOLSTICE to be
re-booted on day 122. Investigation continues into the cause. Otherwise,
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.

***********
SUSIM
***********

SUSIM began a major calibration on May 2.  To ensure accurate solar
scans of its reference channels, closed loop solar pointing of the SSPP
was used.  Lamp calibrations of these reference channels are/were then
performed.  In addition, SUSIM performed its daily solar scans this
past week during periods of predicted UARS telemetry.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

     PEM/AXIS was activated in its full power mode on April 30.
At the current time, AXIS is not generating energy deposition.  
SwRI is looking into this.  Tests of AXIS production code on the rehost 
computer are continuing.

     At the present time, PEM energy deposition from its spectrometers is not 
being generated.  Recent data configuration changes have affected PEM software 
causing multiple failures.  These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update 
processing code with a patch delivery.  The delivery patch may include changes 
which would allow the PEM software to execute on a DEC Alpha machine.  GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).

     PEM has examined its processed files of the rehost machine and has found 
that no differences exist in the Level 1 and Level 2 output files which were 
generated the rehost when the same input data was used on CDHF and CDHFB.
PEM is now confident that output generated by PEM 3.32, Jobs 1 & 2, on the 
rehost system is valid and correct.

     The PEM team is working on replacing the PEM:: machine with a DEC Alpha 
machine to be called PEM2K::.  Tests of the NRT data transfer from CDHF2:: 
through PEM2K:: to the unix computer running the PEM GSE software show no 
difference between equivalent NRT files transferred from PEM::.  Thus, NRT 
data from CDHF2:: is captured correctly on PEM2K::.  Tests of the NRT data 
transfer from CDHFB:: were not successful from both PEM:: and PEM2K::.  
Results of the test of NRT data transfers from the rehost computer were
successful for individual files which exist on CDHFB::.  PEM will participate
in the NRT data flow to capture NRT data from CDHFB:: when specified by GSFC.

     PEM has been working with GSFC on allowing commanding of the PEM
instrument by creating DAPs on the PEM2K machine and submitting them to the 
CMS.  PEM2K:: now is able to communicate with the CMS, but the CMS does
not recognize PEM2K:: as an authorized site for instrument daily activity
plan submission.  PEM is waiting for a solution from GSFC.

    Updated HEPS calibration parameters have been finalized and are ready for 
delivery.  Updated MEPS quality calibration files have been finalized and are 
ready for delivery.  Level 3 software errors, discovered by Dr. Dean Pesnell, 
have been fixed; his suggestions have been incorporated with in the PEM 
production code.  Both updated HEPS and MEPS calibration data, along with the 
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

HALOE was operating successfully with all engineering parameters within 
nominal operating range when it was turned off May 3, 2001 for the solar 
graze period.  All science data products continue to be of excellent 
quality.  HALOE is scheduled to be powered on May 16, 2001.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

The MPG participated in the UARS Monthly Status Review
Meeting on 2 May 2001.  Daily processing of instrument 
Daily Activity Plans continues with no problems. The 
instrument operations schedule is now complete
to 28 June 2001.

***********
CDHF
***********

DAAC DATA
---------
- CLAES (OUTPUT 9) data created in May, 1999 has been previewed in
preparation for sending to the DAAC.  The list of May files has been
forwarded to CLAES for review and approval. 

UARS CDHF REHOST
----------------
- Target date of May 25 has been requested to confirm "puts" to silo.
Once this is done we are able to go operational. We will review NRT user
access status on Rehost next week.

                 
***********
FOT
***********

                        REPORTING PERIOD

This report covers 21 April 2001 (Orbit 52531, GMT Day 111) through 27 April 
2001 (Orbit 52635, GMT Day 117).

                     SPACECRAFT OPERATION

The observatory is now in Forward Flight (Southern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


              INSTRUMENT OPERATIONAL CHANGES


 Instrument     Time       Orbit     Comment          
 ACRIM        112/0054     52546     Power ON
 SUSIM        112/0316     52547     Power ON
 SUSIM        112/1622     52556     Power ON complete
 HALOE        113/2357     52575     Power ON (Part 1)
 HALOE        114/0457     52578     Power ON (complete)          

The beta angle ranged between the angles of 5.6 degrees to 31.6 degrees for 
this report period.  The beta angle is now increasing towards a maximum angle 
of 74 degrees on 09 May 2001 (DOY 129).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 1.5 and 1.6 degrees C.  Additional battery performance data for each 
battery on 20 April 2001 and 27 April 2001 are provided in Appendix D.

The clock error ranged from 12.4 to 10.4 msec during this report period.  
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

     OUT-OF-LIMITS  =  02
     ALERTS         =  00

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).

     FOT            =  17
     HALOE          =  02
     SUSIM          =  02
     ACRIM          =  01
     PEM            =  01

UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.

     FOT           =  01

One (1) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/316  - 2287.496493
 Transponder B:  01/087  - 2287.499152


                    GROUND SYSTEM OPERATION

One (1) operational support problems occurred this period.  Anomaly Reports are 
included in Appendix E.

The data loss calculations supplied by Data Capture Facility (DCF) from mission 
start through UTC Day 01/067 show the observatory data loss to be 300 hours, 18 
minutes, 29.768 seconds. This shows an increase of 15 minutes and 24.304 
seconds since the last report period. This is a 0.39 percent data loss, which 
equals a 99.61 percent data capture for the mission.

The increase of 15 minutes and 24.304 seconds occurred over four (4) days:
     1 minute and 6.384 seconds on Day 110 due to a Data Hit,
     1 minute and 16.368 seconds on Day 111 due to a Data Hit.
     10 minutes and 47.536 seconds on Day 112 with reason unknown.
     2 minutes and 14.016 seconds on Day 112 with reason unknown.

                   OTHER SIGNIFICANT EVENTS

                   UARS SATELLITE OPERATIONS

The new flight software version 67 was successfully uplinked on 25 April 2001.  
This patch increases the accuracy of the magnetic field model, accounts for the 
one second delay between MAGMOD and UFLTR execution and properly calculates the 
magnetic field model time.  These enhancements should provide for more accurate 
TAM measurements and increase the overall attitude precision. 

After the uplink the FOT also switched from using TAM-2 to using TAM-1.  The 
switch was made due to the fact that TAM-1 is more orthogonal than TAM-2.  The 
TAM-1 alignment will be fine-tuned by loading new alignment matrices this week.
An accurate alignment is needed to get ready for the loss of sun in the field 
of view of the FSS on 5 May 2001.  


                           MEETINGS

The FOT attended a UARS Flight Code Review Meeting on Monday, April 23, 2001.
The FOT attended a Hazard Training Class on Wednesday, April 25, 2001.

                    FUTURE SIGNIFICANT EVENTS

The next UARS Yaw slew is scheduled for 30 May 2001 (DOY 150).  This maneuver 
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).


Attachments:  Appendix A,  Real-time Commands Sheets
              Appendix B,  Out-Of-Limits & ALERT Occurrences
              Appendix C,  Clock & Solar Array Rate Corrections
              Appendix D,  UARS Battery Performance Data
              Appendix E,  UARS Anomaly Reports

                         APPENDIX A
                 REAL-TIME COMMAND SHEETS
                 Week ending 27 April 2001

ORBIT     AOS         NO.      SUBSYS     REASON
-----     ---         ---      ------     ------
52531     111/0136    01-253   OBC        Clk Adj 0.0 to -20.6 msec/day
52546     112/0054    01-254   SOLSTICE   From Day-Only to Fulltime Ops
52546     112/0054    01-255   ACRIM      Power ON
52547     112/0316    01-256   SUSIM      Power ON
52549     112/0640    01-257   OBC        Clk Adj -20.6 to 0.0 msec/day
52553     112/258     01-258   PEM        ZEPS & HEPS to Day-Only ops.
52556     112/1622    01-256   SUSIM      Power ON complete
52572     113/1812    01-259   NBTR       NOOPS D-114
52575     113/2357    01-261   HALOE      Power ON (Part 1)
52578     114/0457    01-262   HALOE      Power ON (complete)
52584     114/1337    01-259   NBTR       NOOPS D-114
52587     114/1813    01-263   NBTR       NOOPS D-115
52599     115/1439    01-264   OBC        FSW ver# 67 update
52600     115/1518    01-265   OBC        TAM 2 TO TAM 1
52601     115/1628    01-266   OBC        TAM to Update Mode
52602     115/1815    01-267   ATC        NOOPS D-116
52610     116/0756    01-267   NBTR       NOOPS D-116
52616     116/1704    01-267   NBTR       NOOPS D-116
52617     116/1816    01-268   NBTR       NOOPS D-117
52623     117/0508    01-268   NBTR       NOOPS D-117
52624     117/0616    01-268   NBTR       NBTP PBK / RTS-58
52630     117/1522    01-269   OBC        Clk Adj 0.00 to 20.6 msec/day
52631     117/1633    01-268   NBTR       NOOPS D-117
52633     117/2038    01-270   NBTR       NOOPS D-118 
________________________________________________________________________

                            APPENDIX B
            OUT-OF-LIMITS & ALERTS OCCURRENCES
                 Week ending 27 April 2001

The following Out-Of-Limits occurrences were detected:

MNEMONIC       STATE     DESCRIPTION                   ORBITS
--------       -----     -----------                    ------
HRIFESURVTMP   Y-HI      HRDI PSE Quiet Bus Cur Mon     52611
HRRADTEMP      Y-HI      HRDI PSE Pulse Bus Cur Mon     52611


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION     ORBITS
--------     -----------     ------
-None this report period

_______________________________________________________________________________

                        APPENDIX C
           CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 April 2001

CLOCK DRIFT RATE ADJUSTMENTS

ORBIT     TIME     CLOCK RATE CHANGE
-----     -----     -----------------

52531     111/0141     Clk Adj 0.0 to -20.6 msec/day
52549     112/0643     Clk Adj -20.6 to 0.0 msec/day
52630     117/1524     Clk Adj 0.00 to 20.6 msec/day
___________________________________________________________________________

                        APPENDIX D
             UARS BATTERY PERFORMANCE DATA
               Week ending 27 April 2001


     20 Apr 2001 - GMT Day 110 Beta = 01.4 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)      CURR (amp)    EON V    AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN       MAX/MIN       MIN      RATIO     DOD(%)
---  ----------     -------       ---------     ------   -------   ------
1    +621.6/+280.0  -3.68/-4.94   +0.0/+0.0     12.3     0.000     0.0
2    +22.4/-67.2    +5.15/+3.87   +28.8/-12.8   25.1     1.020     20.5
3    +22.4/-67.2    +3.87/+2.60   +28.0/-12.8   25.0     1.020     20.0


     27 Apr 2001 - GMT Day 117 Beta = 31.6 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)      CURR (amp)    EON V    AVE C/D    AVE
BAT  MAX/MIN        MAX/MIN       MAX/MIN       MIN      RATIO      DOD(%)
---  ----------     -------       ---------     ------   -------    ------
1    +728.0/+728.0  -3.05/-4.31   +0.0/+0.0     12.0     0.000      0.0
2    +11.2/-89.6    +6.11/+4.51   +28.8/-15.2   24.2     1.030     21.6
3    +22.4/-50.4    +5.15/+3.55   +28.8/-16.0   24.2     1.020     21.6

     V/T LEVEL CHANGES
ORB#             TIME          LVL FR-TO
----             ----          ---------
-no changes      V/T 5

_____________________________________________________________________________

                              APPENDIX E
                   UARS ANOMALY REPORTS
                  Week ending 27 April 2001

                NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
01-020     FOT        52609     116/0615   N/A     TPF Late ACQ on PBK

           ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
01-020     FOT        52609     116/0615   N/A     TPF Late ACQ on PBK

            ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
-None this report period

                       OPEN ANOMALY REPORTS
AR NO.     SUBSYS     ORBIT     TIME       TTR     PROBLEM/DATA LOSS
------     ------     -----     ----       ---     -----------------
01-019     FOT        see list  Apr 01     14170   Generic Late Acquisition


	 	SEVNTFW01/16